Please find thickness

Pressure (PSI) Hoop Stress Thickness in. Longitudinal Stress Thickness in. 500 1000 1500 1000 500 Average Thickness Average Thickness: in.
Answer #1

\small \sigmah = pd/2t

for Design, \small \sigmah\tiny \geq pd/2t

t \tiny \geq pd / 2\small \sigmah as per hoop stress

Similarly as per longitudinal stress, t \tiny \geq pd / 4\small \sigmaL

Diameter, d = 2 in

Pressure (PSI)

Hoop Stress

(\small \sigmah ) (PSI)

Hoop Stress Thickness

(th) (in)

Longitudinal Stress

(\small \sigmaL ) (PSI)

Longitudinal Stress

Thickness (tL) (in)

500 6031.41 0.083 2897.54 0.086
1000 11426.44 0.088 5439.35 0.092
1500 17880.05 0.084 8525.21 0.088
1000 1948.34 0.513 5680.60 0.088
500 6100.34 0.082 2493.08 0.100
Average Thickness 0.170 0.091

As per hoop stress, thickness t = 0.17 in

As per longitudinal stress, thickness t = 0.091

Hence Average Thickness, t =0.131 in